One angle of a triangle is twice as large as another and 25 degree more than the third angle. Find the angles.
Answer by sachi(548)

let the 1st angle is x
2nd angle is x/2
3rd angle is x-25
now x+x/2+x-25=180
or 5x/2=180+25=205
or x=205*2/5=82
so the angles are 82,41,57 ans